# Eulerian magnitude homology: diagonality, injective words, and regular path homology

### v1

@@ 2. Regularly diagonal graphs
[…]
@@
If $G$ is regularly diagonal, then $\mathrm{EMH}_{k,\ell}(G)$ is trivial for every $k\neq\ell$, and substituting the trivial values of eulerian magnitude homology in the long exact sequence in Equation 3 the result follows. ∎
@@ 3. Regularly diagonal directed graphs
We have seen that eulerian magnitude homology of undirected graphs detects the complete graphs. The story is completely different when we consider the bigger family directed graphs. In this section we show that taking cones and joins preserves diagonality, and as an example we compute the (eulerian) magnitude homology of transitive tournaments.
The simplest family of directed graphs which are regularly diagonal, but not of the form $\rho(K_{n})$, is the family of alternating digraphs. For an undirected graph $G$ and an orientation $o$ of its edges, we say that $o$ is alternating if there exists a partition $V\sqcup W$ of the vertices of $G$ such that all elements of $V$ have indegree 00 and all elements of $W$ have outdegree 00cf[CCC24]. The directed graph $(G,o)$ is the undirected graph $G$ equipped with the orientation $o$ of its edges. Then, a direct computation shows the following:
@@
$\mathrm{rk}\,\mathrm{MH}_{k,l}(G,o)=\mathrm{rk}\,\mathrm{EMH}_{k,l}(G,o)=\begin{cases}|V(G)|&\text{ if }(k,l)=(0,0)\\ |E(G)|&\text{ if }(k,l)=(1,1)\\ 0&\text{ otherwise }\end{cases}$
@@
Let $G$ be a connected graph. Notice that it is possible to construct the directed graph $(G,o)$ if and only if $G$ is bipartite. Indeed, if $G=(V,E)$ is bipartite then $V=V_{1}\cup V_{2}$ with $V_{1}\cap V_{2}=\emptyset$, and it is possible to choose an orientation $o$ of the edges such that all vertices in $V_{1}$ have indegree 00 and all elements of $V_{2}$ have outdegree 00. On the other hand, say we have a directed graph $(G,o)$ and suppose $G$ is not bipartite. Then it not possible to divide the vertex set $V$ of $G$ in two disjoint independent sets, meaning for every choice of $V_{1},V_{2}$ such that $V_{1}\cup V_{2}=V$ there exists a vertex $v\in V_{1}\cap V_{2}$. But then by construction of $(G,o)$ the vertex $v$ would have both indegree and outdegree equal to 00, contradicting the connectivity of $G$.
@@
where the sum is taken across all vertices $x_{1}<x_{k}$ of $L_{n}$ at distance $\ell$; here $\mathrm{EMH}_{*,\ell}(x_{1},x_{k})$ denotes the sub-chain complex consisting of all chains starting at $x_{1}$ and ending at $x_{k}$. Now, observe that each such term $\mathrm{EMH}_{*,\ell}(x_{1},x_{k})$ is the homology of a Boolean poset, which is trivial. As a consequence, for any $\ell>1$, the (eulerian) magnitude homology groups $\mathrm{EMH}_{k,\ell}(L_{n})$ are trivial. It is left to consider the case of bidegrees $(0,0)$ and $(1,1)$. In such cases, the (eulerian) magnitude homology groups are of rank $n$ and $n-1$, respectively.
@@
The class of bipartite graphs and of trees provides many non-trivial examples of regularly diagonal digraphs. We shall also construct regularly diagonal graphs by means of taking cones and joins. First recall that the cone $C(G)$ of a digraph $G$ is defined as follows. The vertices of $C(G)$ are the vertices of $G$ with an extra vertex $\star$ added. The directed edges of $C(G)$ are the directed edges of $G$, together with all directed edges of type $(x,\star)$ for all $x$ in $G$.
@@
$\mathrm{EMH}_{k,\ell}(C(G))\cong\begin{cases}\mathrm{EMH}_{k,\ell}(G)\oplus\mathbb{Z}&\text{ if }(k,\ell)=(0,0);\\ \mathrm{EMH}_{k,\ell}(G)\oplus\mathrm{EMH}_{k-1,\ell-1}(G)&\text{ otherwise }.\end{cases}$
@@
of modules: an element of $\mathrm{EMC}_{k,\ell}(C(G))$ either corresponds to a tuple $(x_{0},\dots,x_{k})$ in $G$, or to a tuple of type $(x_{0},\dots,x_{k-1},\star)$. The differential on the first component is the differential of $\mathrm{EMC}_{*,*}(G)$. Consider the second component. In such case, we have $\partial_{i}(x_{0},\dots,x_{k-1},\star)=(x_{0},\dots,\hat{x_{i}},\dots,x_{k-1},\star)$ for all $i=0,\dots,k_{1}$. The last face map $\partial_{k}$ sends $(x_{0},\dots,x_{k-1},\star)$ to 00, as $(x_{0},\dots,x_{k-1})$ is of length $\mathrm{len}(x_{0},\dots,x_{k-1},\star)-1$. Hence, also the differential on $\mathrm{EMC}_{k,\ell}(C(G))$ is consistent with the differential on the decomposition $\mathrm{EMC}_{k,\ell}(G)\oplus\mathrm{EMC}_{k-1,\ell-1}(G)$; showing that Eq. 4 is in fact an isomorphism of chain complexes. This is enough to show that $\mathrm{EMH}_{k,\ell}(C(G)$ is isomorphic to $\mathrm{EMH}_{k,\ell}(G)\bigoplus\mathrm{EMH}_{k-1,\ell-1}(G)$. ∎
@@
We start with computing the (eulerian) magnitude homology groups of $T_{n}$ in bidegree $(2,2)$. Although there are no cycles of length $3$, for each pair of vertices $i,j$ with $i<j$, there are $j-i-1$ paths of length $2$ from $i$ to $j$, meaning $\mathrm{rk}\,\mathrm{EMH}_{2,2}(i,j)=j-i-1$. Here $\mathrm{EMH}_{2,2}(i,j)$ indicates the subgroup of $\mathrm{EMH}_{2,2}(G)$ generated by tuples starting at $i$ and ending at $j$. Therefore, accounting for all possible pairs $(i,j)$ with $i<j$, for $n\geq 2$ we have
@@
based on the fact that to compute the number of ordered increasing sequences of $k$-tuples it is enough to fix the last (or initial) digits and then to compute the number of $(k-1)$-tuples on the resulting subsequences. Putting together, we have that the rank of $\mathrm{EMH}_{k,k}(T_{n})$ equals the sum:
@@
More generally, we can extend the result of Proposition 3.5 to the directed joins. We define the directed join of the directed graphs $G$ and $H$ to be the digraph $G*H$ on the vertices $V(G)\sqcup V(H)$ with edges
@@
For $G*H$ directed join of $G$ and $H$, the following formula holds:
@@
First, observe that the eulerian magnitude chains of $G*H$ consist of tuples $(x_{0},\dots,x_{k_{1}},y_{0},\dots,y_{k_{2}})$, with $(x_{0},\dots,x_{k_{1}})\in\mathrm{EMC}_{k_{1},l_{1}}$ and $(y_{0},\dots,y_{k_{2}})\in\mathrm{EMC}_{k_{2},l_{2}}$. As each vertex of $G$ is connected to each vertex of $H$ by a directed edge, the length of $(x_{0},\dots,x_{k_{1}},y_{0},\dots,y_{k_{2}})$, is $l=l_{1}+l_{2}+1$. Hence, as modules, we have a decomposition of $\mathrm{EMC}_{k,l}(G*H)$ as direct sum $\mathrm{EMC}_{k_{1},l_{1}}(G)\oplus\mathrm{EMC}_{k_{2},l_{2}}(H)$, via the map
@@ 4. The complex of injective words on a digraph
In this section, when otherwise stated, $G$ shall denote a directed graph without multiple directed edges.
Recall the definition of complex of injective words on a finite set $W$:
@@
The complex of injective words $\Delta(W)$ on the set $W$ is the simplicial complex on the injective words words $w=v_{1}v_{2}\cdots v_{k}$ of $W$ with alphabet the elements of $W$. The $p$-simplices are the injective words with $p+1$ letters.
Saying otherwise, the simplicial complex $\Delta(W)$ is the simplicial complex on all ordered sequences of distinct elements of $W$. For $n\geq 1$, we let $D(n)\coloneqq\sum_{j=0}^{n}(-1)^{j}\frac{n!}{j!}$ denote the number of derangements (that is, the fixed point free permutations) in the symmetric group $\Sigma_{n}$. The following theorem shows that the homotopy type of the complex of injective words is quite simple:
@@
We now wish to extend the notion of complexes of injective words to directed graphs.
@@
For a directed graph $G$ with vertex set $V$, we define the complex of injective words on $G$ to be the simplicial complex $\mathrm{Inj}(G)$ generated by all injective words $w=x_{0}\cdots x_{k}$ on the alphabet $V$, with the requirement that for all $i<j$ there is a directed path in $G$ from $x_{i}$ to $x_{j}$.
@@
Using Notation 1.1, if $G=\rho(K_{n})$ is the complete digraph with both reciprocal edges between any pair of vertices, then $\mathrm{Inj}(G)=\Delta(V)$, the classical complex of injective words. In the general case, we shall see that $\mathrm{Inj}(G)$ is a directed flag complex, hence the complexity of its homotopy type can be arbitrary. First, consider the transitive closure of $G$. Note that the transitive closure yields a preorder, since the graph $G$ has no multiple directed edges. The vertices of the transitive closure of $G$ are the vertices of $G$ and there is a directed edge $(x,y)$ if and only if there is a path from $x$ to $y$. Note that, when related to magnitude homology, the transitive closure of $G$ is also called the reachability preorder $\mathrm{Pre}(G)$ of $G$cf[DIMZ24, CR24, HR23]. We will also follow this notation. Passing to the reachability preorder does not affect the complex of injective words:
Lemma 4.5.
@@
We wish to interpret the complex of injective words on a directed graphs in terms of directed flag complexes. Recall first that an ordered $k$-clique of a directed graph $G$ is a totally ordered $k$-tuple $(v_{1},...,v_{k})$ of vertices of $G$ with the property that, for every $i<j$, the pair $(v_{i},v_{j})$ is an ordered edge of $G$. Then, for a given directed graph $G$, we can construct the associated directed flag complex $\mathrm{dFl}(G)$, that is the ordered simplicial complex on the vertices $V$ of $G$ whose $k$-simplices are all possible ordered $(k+1)$-cliques of $G$.
@@
Let $G$ be a digraph. Then, $\mathrm{Inj}(G)\cong\mathrm{dFl}(\mathrm{Pre}(G))$.
@@
sending an injective word $x_{0}\cdots x_{k}$ to the clique $(x_{0},\dots,x_{k})$. The map is well-defined because, by definition, for all $i<j$ the distinct vertices $x_{i}$ and $x_{j}$ are connected by a directed path in $G$ (hence, by an edge in $\mathrm{Pre}(G)$), and they yield a clique in $\mathrm{Pre}(G)$. Both the differentials on $\mathrm{Inj}(G)$ and $\mathrm{dFl}(\mathrm{Pre}(G))$ act by forgetting the $i$-th entry, hence the map is simplicial. Furthermore, it is an isomorphism of simplicial complexes with inverse the map $\mathrm{dFl}(\mathrm{Pre}(G))\to\mathrm{Inj}(G)$ sending the clique $(x_{0},\dots,x_{k})$ to the injective word $x_{0}\cdots x_{k}$. ∎
Let $H$ be an undirected graph on $n$ vertices. If $G=\rho(H)$, then its reachability preorder $\mathrm{Pre}(G)$ is the complete digraph $\rho(K_{n})$ on the same number of vertices. Then, the associated directed flag complex $\mathrm{dFl}(\mathrm{Pre}(G))$ is isomorphic to $\mathrm{dFl}(\rho(K_{n}))=\Delta(V(G))$, finding again the standard complex of injective words.
@@
If $G$ is the directed graph associated to a poset $P$, that is the graph on the same vertices of $P$ with an edge $(x,y)$ if and only if $x\leq y$ in $P$, then $\mathrm{dFl}(\mathrm{Pre}(G))$ is the standard order complex of $P$ [^1] Recall that the order complex of a poset $(P,\leq)$ is the simplicial complex on the vertex set $P$ whose $k$-simplices are the chains $x_{0}<\cdots<x_{k}$ of $P$..
By Theorem 4.2, the complex of injective words is always a wedge of spheres. This is not the case for the complex of injective words on directed graphs, which can have any homotopy type. In fact, as the transitive closure $\mathrm{Pre}(P)$ of a poset $P$ is the poset $P$ itself, we have $\mathrm{Inj}(P)\cong\mathrm{dFl}(P)$, that is the order complex of $P$.
For any simplicial complex $\Sigma$ we can consider its face poset $F(\Sigma)$, whose nerve is (the barycentric subdivision of) $\Sigma$. Then, $\mathrm{Inj}(F(\Sigma))$ is homotopy equivalent to $\Sigma$, and the complex of injective words on directed graphs has any homotopy type.

### v2

@@ 2. Regularly diagonal graphs
$0\to\mathrm{EMC}_{\ast,\ell}(G)\xrightarrow{\iota}\mathrm{MC}_{\ast,\ell}(G)\xrightarrow{\pi}\mathrm{DMC}_{\ast,\ell}(G)\to 0\ ,$
@@
If $G$ is regularly diagonal, then $\mathrm{EMH}_{k,\ell}(G)$ is trivial for every $k\neq\ell$. As $\mathrm{DMH}_{\ell,\ell}(G)$ is free, the result follows. ∎
@@ 3. Regularly diagonal directed graphs
We have seen that eulerian magnitude homology of undirected graphs detects the complete graphs. The story is completely different when we consider the bigger family of directed graphs. In this section we show that taking cones and joins preserves diagonality, and as an example we compute the (eulerian) magnitude homology of transitive tournaments.
First, observe that the simplest family of directed graphs which are regularly diagonal, but not of the form $\rho(K_{n})$, is the family of alternating digraphs. For an undirected graph $G$ and an orientation $o$ of its edges, we say that $o$ is alternating if there exists a partition $V\sqcup W$ of the vertices of $G$ such that all elements of $V$ have indegree 0 and all elements of $W$ have outdegree 0cf[CCC24]. The directed graph $(G,o)$ is the undirected graph $G$ equipped with the orientation $o$ of its edges. Then, a direct computation shows the following:
@@
[…]
@@
Let $G$ be a connected graph. Notice that it is possible to construct the directed graph $(G,o)$ if and only if $G$ is bipartite. Indeed, if $G=(V,E)$ is bipartite then $V=V_{1}\cup V_{2}$ with $V_{1}\cap V_{2}=\emptyset$, and it is possible to choose an orientation $o$ of the edges such that all vertices in $V_{1}$ have indegree 0 and all elements of $V_{2}$ have outdegree 0. On the other hand, say we have a directed graph $(G,o)$ and suppose $G$ is not bipartite. Then it not possible to divide the vertex set $V$ of $G$ in two disjoint independent sets, meaning for every choice of $V_{1},V_{2}$ such that $V_{1}\cup V_{2}=V$ there exists a vertex $v\in V_{1}\cap V_{2}$. But then by construction of $(G,o)$ the vertex $v$ would have both indegree and outdegree equal to 0, contradicting the connectivity of $G$.
@@
where the sum is taken across all vertices $x_{1}<x_{k}$ of $L_{n}$ at distance $\ell$; here $\mathrm{EMH}_{*,\ell}(x_{1},x_{k})$ denotes the sub-chain complex consisting of all chains starting at $x_{1}$ and ending at $x_{k}$. Now, observe that each such term $\mathrm{EMH}_{*,\ell}(x_{1},x_{k})$ is the homology of a Boolean poset, which is trivial. As a consequence, for any $\ell>1$, the (eulerian) magnitude homology groups $\mathrm{EMH}_{k,\ell}(L_{n})$ are trivial. It is left to consider the case of bidegrees $(0,0)$ and $(1,1)$. In such cases, the (eulerian) magnitude homology groups are of rank $n$ and n1n-1, respectively.
@@
The class of bipartite graphs and trees provides many non-trivial examples of regularly diagonal digraphs. We shall also construct regularly diagonal graphs by means of taking cones and joins. First, recall that the cone $C(G)$ of a digraph $G$ is defined as follows. The vertices of $C(G)$ are the vertices of $G$ with an extra vertex $\star$ added. The directed edges of $C(G)$ are the directed edges of $G$, together with all directed edges of type $(x,\star)$ for all $x$ in $G$.
@@
[…]
@@
of modules: an element of $\mathrm{EMC}_{k,\ell}(C(G))$ either corresponds to a tuple $(x_{0},\dots,x_{k})$ in $G$, or to a tuple of type $(x_{0},\dots,x_{k-1},\star)$. The differential on the first component is the differential of $\mathrm{EMC}_{*,*}(G)$. Consider the second component. In such case, we have $\partial_{i}(x_{0},\dots,x_{k-1},\star)=(x_{0},\dots,\hat{x_{i}},\dots,x_{k-1},\star)$ for all $i=0,\dots,k_{1}$. The last face map $\partial_{k}$ sends $(x_{0},\dots,x_{k-1},\star)$ to 0, as $(x_{0},\dots,x_{k-1})$ is of length $\mathrm{len}(x_{0},\dots,x_{k-1},\star)-1$. Hence, also the differential on $\mathrm{EMC}_{k,\ell}(C(G))$ is consistent with the differential on the decomposition $\mathrm{EMC}_{k,\ell}(G)\oplus\mathrm{EMC}_{k-1,\ell-1}(G)$; showing that Eq. 4 is in fact an isomorphism of chain complexes. This is enough to show that $\mathrm{EMH}_{k,\ell}(C(G)$ is isomorphic to $\mathrm{EMH}_{k,\ell}(G)\bigoplus\mathrm{EMH}_{k-1,\ell-1}(G)$. ∎
@@
We start with computing the (eulerian) magnitude homology groups of $T_{n}$ in bidegree $(2,2)$. Although there are no cycles of length $3$, for each pair of vertices $i,j$ with $i<j$, there are ji1j-i-1 paths of length $2$ from $i$ to $j$, meaning $\mathrm{rk}\,\mathrm{EMH}_{2,2}(i,j)=j-i-1$. Here $\mathrm{EMH}_{2,2}(i,j)$ indicates the subgroup of $\mathrm{EMH}_{2,2}(G)$ generated by tuples starting at $i$ and ending at $j$. Therefore, accounting for all possible pairs $(i,j)$ with $i<j$, for $n\geq 2$ we have
@@
based on the fact that to compute the number of ordered increasing sequences of $k$-tuples it is enough to fix the last (or initial) digits and then to compute the number of (k1)(k-1)-tuples on the resulting subsequences. Putting together, we have that the rank of $\mathrm{EMH}_{k,k}(T_{n})$ equals the sum:
@@
More generally, we can extend the result of Proposition 3.5 to the directed joins. We define the directed join of the directed graphs $G$ and $H$ to be the digraph GHG*H on the vertices $V(G)\sqcup V(H)$ with edges
@@
For GHG*H directed join of $G$ and $H$, the following formula holds:
@@
First, observe that the eulerian magnitude chains of GHG*H consist of tuples $(x_{0},\dots,x_{k_{1}},y_{0},\dots,y_{k_{2}})$, with $(x_{0},\dots,x_{k_{1}})\in\mathrm{EMC}_{k_{1},l_{1}}$ and $(y_{0},\dots,y_{k_{2}})\in\mathrm{EMC}_{k_{2},l_{2}}$. As each vertex of $G$ is connected to each vertex of $H$ by a directed edge, the length of $(x_{0},\dots,x_{k_{1}},y_{0},\dots,y_{k_{2}})$, is $l=l_{1}+l_{2}+1$. Hence, as modules, we have a decomposition of $\mathrm{EMC}_{k,l}(G*H)$ as direct sum $\mathrm{EMC}_{k_{1},l_{1}}(G)\oplus\mathrm{EMC}_{k_{2},l_{2}}(H)$, via the map
@@ 4. The complex of injective words on a digraph
In this section, we first recall the complex of injective words on a set, and extend this notion to directed graphs.
Let $W$ be a finite set.
@@
The complex of injective words $\Delta(W)$ on the set $W$ is the simplicial complex on the injective words $w=v_{1}v_{2}\cdots v_{k}$ of $W$ with alphabet the elements of $W$. The $p$-simplices are the injective words with $p+1$ letters.
Saying otherwise, the simplicial complex $\Delta(W)$ is the simplicial complex on all ordered sequences of distinct elements of $W$. For $n\geq 1$, we let $D(n)\coloneqq\sum_{j=0}^{n}(-1)^{j}\frac{n!}{j!}$ denote the number of derangements (that is, the fixed point free permutations) in the symmetric group $\Sigma_{n}$. The following theorem shows that the homotopy type of the complex of injective words is well-understood:
@@
We now extend the notion of complexes of injective words to directed graphs.
@@
For a directed graph $G$ with vertex set $V(G)$, we define the complex of injective words on $G$ to be the simplicial complex $\mathrm{Inj}(G)$ generated by all injective words $w=x_{0}\cdots x_{k}$ on the alphabet $V(G)$, with the requirement that for all $i<j$ there is a directed path in $G$ from $x_{i}$ to $x_{j}$.
@@
Example 4.5.
If $BK_{n}=\rho(K_{n})$ is the complete digraph with both reciprocal edges between any pair of vertices, then $\mathrm{Inj}(BK_{n})=\Delta(V)$, the classical complex of injective words on the set $V$ of vertices of $BK_{n}$.
In the general case, we shall see that $\mathrm{Inj}(G)$ is a directed flag complex, hence the complexity of its homotopy type can be arbitrary. To see it, consider the transitive closure of $G$. Its vertices are the vertices of $G$ and there is a directed edge $(x,y)$ if and only if there is a directed path from $x$ to $y$ – including the trivial constant path. The transitive closure yields a preorder, since the graph $G$ has no multiple directed edges. In the literature related to magnitude homology, this preorder is also called the reachability preorder $\mathrm{Pre}(G)$ of GGcf[DIMZ24, CR24, HR23]. We will also follow this notation. First, observe that passing to the reachability preorder does not affect the complex of injective words:
Remark 4.6.
@@
Recall that an ordered $k$-clique of a directed graph $G$ is a totally ordered $k$-tuple $(v_{1},...,v_{k})$ of vertices of $G$ with the property that, for every $i<j$, the pair $(v_{i},v_{j})$ is an ordered edge of $G$. Then, for a given digraph $G$, we can construct the associated directed flag complex $\mathrm{dFl}(G)$, that is the ordered simplicial complex on the vertices $V(G)$ of $G$ whose $k$-simplices are all possible ordered $(k+1)$-cliques of $G$.
@@
Let $G$ be a digraph. Then, we have an isomorphism $\mathrm{Inj}(G)\cong\mathrm{dFl}(\mathrm{Pre}(G))$.
@@
sending an injective word $x_{0}\cdots x_{k}$ to the ordered clique $(x_{0},\dots,x_{k})$. The map is well-defined because, by definition, for all $i<j$ the distinct vertices $x_{i}$ and $x_{j}$ are connected by a directed path in $G$ (hence, by an edge in $\mathrm{Pre}(G)$), and they yield an ordered clique in $\mathrm{Pre}(G)$. Both the differentials on $\mathrm{Inj}(G)$ and $\mathrm{dFl}(\mathrm{Pre}(G))$ act by forgetting the $i$-th entry, hence the map is simplicial. Furthermore, it is an isomorphism of simplicial complexes with inverse the map $\mathrm{dFl}(\mathrm{Pre}(G))\to\mathrm{Inj}(G)$ sending the ordered clique $(x_{0},\dots,x_{k})$ to the injective word $x_{0}\cdots x_{k}$. ∎
Let $H$ be an undirected graph on $n$ vertices, and $K_{n}$ be the complete graph on $n$ vertices. Consider $G=\rho(H)$. Then, the reachability preorder $\mathrm{Pre}(G)$ is isomorphic to the complete digraph $\rho(K_{n})$. Consequently, the associated directed flag complex $\mathrm{dFl}(\mathrm{Pre}(G))$ is isomorphic to $\mathrm{dFl}(\rho(K_{n}))=\Delta(V(G))$, the standard complex of injective words.
@@
If $G$ is the directed graph associated to a poset $P$, that is the graph on the same vertices of $P$ with an edge $(x,y)$ if and only if $x\leq y$ in $P$, then $\mathrm{dFl}(\mathrm{Pre}(G))$ is the standard order complex of PP [^1] Recall that the order complex of a poset $(P,\leq)$ is the simplicial complex on the vertex set $P$ whose $k$-simplices are the chains $x_{0}<\cdots<x_{k}$ of $P$..
By Theorem 4.2, the complex of injective words is, up to homotopy, always a wedge of spheres. This is not the case for the complex of injective words on digraphs, which can have any homotopy type. To see it, first observe that as $\mathrm{Pre}(P)\cong P$, we have an isomorphism $\mathrm{Inj}(P)\cong\mathrm{dFl}(P)$ – that is the order complex of $P$. Given a simplicial complex $\Sigma$, consider its face poset $P=F(\Sigma)$, whose nerve is (the barycentric subdivision of) $\Sigma$. Then, the complex of injective words $\mathrm{Inj}(F(\Sigma))$ is homotopy equivalent to $\Sigma$; which was arbitrary and can have any homotopy type.